Circular No. 5608

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

AM HERCULIS N. M. Shakhovskoj, S. V. Kolesnikov, I. L. Andronov, and I. Yu. Alex-

eev write: “An unprecedented flare of AM Her during its recent deep low state was observed simultaneously with the 2.6-m and 1.25-m telescopes of the Crimean Astrophysical Observatory. After a phase of nearly constant brightness (V = 15.40, U–B = –0.68, B–V = +0.18, V–R = +0.61, R–I = +1.23; time resolution 24 s at the 1.25-m telescope), a flare to V = 13.26 (U–B = –0.85, B–V = –0.04, V–R = +0.18, R–I = +0.36) occurred on Aug. 29.86123 UT. The rise of 2.12 mag in R (time resolution 4 s, 2.6-m telescope) lasted 2.4 min. The mean value of circular polarization during the flare was zero within an error of 0.1 percent. Characteristic e-folding time was 6.1 min in B. The amplitudes, colors, and timescale argue for a flare of a secondary M-dwarf in a binary system, rather than for an accretion event with cyclotron emission.”

GRO J0422+32 A. Vikhlinin, A. Finoguenov, A. Sitdikov, R. Sunyaev, A. Goldwurm,

J. Paul, P. Mandrou, and P. Techine, on behalf of the Granat team (Space Research Institute, Moscow; Service d’Astrophysique du CEA, Saclay; Centre d’Etude Spatiale de Rayonnements, Toulouse), report: “GRO J0422+32 has been observed by SIGMA/Granat on Aug. 22 in high-time-resolution mode. The quasiperiodic oscillations of flux 40-150 keV were clearly detected at two different frequencies. The QPO peak centroid at 0.28 Hz has FWHM = 0.13 Hz and is bright in the whole 40- to 150-keV band. The QPO peak centered at 0.035 Hz (FWHM = 0.02 Hz) is bright mainly in the 40- to 70-keV band. QPO detection for this source was reported also by BATSE (IAUC 5592). During subsequent observations of GRO J0422+32 from Aug. 15 to Sept. 6, the 40- to 150-keV source flux was decreasing monotonically from 2.91 ± 0.01 Crab to 1.85 ± 0.01 Crab, with an approximate exponential decay time of 44 ± 2 days. This result is based on 270 hr of observation, divided into 12 separate sessions.”

NOVA CYGNI 1992 AND NOVA SCORPII 1992 Corrigendum. The CCD magnitudes from W. Liller reported on IAUC

5605 were of N Sco 1992, not N Cyg 1992.

1992 September 9 Daniel W. E. Green